Calculate the mean free path of H2
molecules at 300 K. The diameter of H2
molecules is 3 Å and the number density is
3 x 10^25 m^-3
1
Expert's answer
2019-05-21T08:43:49-0400

mean free path

"\bar{\lambda}=\frac{1}{\sqrt{2}\pi\sigma^2n}"

calculating

"\bar{\lambda}=\frac{1}{\sqrt{2}\pi(3\cdot 10^{-10})^2\cdot 3\cdot 10^{25}}=83.4\cdot 10^{-9} (m)"

"\bar{\lambda}=83.4 (nm)"